Minutes of nanoBPM meeting on 14 October, 2004

"Status of the system", Y.Honda

All the equipments were installed except for two sets of laser sources and fibers which will be delivered at the end of this October, where total 4 sets are necessary for the laser interferometer and position monitors in the three mover systems. The mover systems are controlled by a PC and data will be read out by another PC. So, we have 2 PC's in our system. At present, we could use spared PC of the laser-wire system. However, a new PC is preferable. In this week, we will check signals of the BPMs without requesting a dedicated machine time.

Two sets of improved electronics will be delivered soon. After we confirm the improvement, the first electronics will be updated.

"Status of KEK nano-BPM", Y. Inoue

The electrical center of the cavity-BPM was measured by the antenna method. Before the measurement, alignment of the antenna was checked by the goniometer and the rotational table. Th results show that the antenna was aligned within -6.6 μrad and -20μrad, respectively. The electrical center was measured to be shifted to -50μm and -140μm in X and Y direction, respectively, from the mechanical one. Although this measurement was not completely understood, we will install the BPMs after the meeting. The electrical center will be checked by beam.

After the meeting, Hayano reported as follows. We installed three cavity BPMs in the line. Please look at pictures (pdf, 3 pages, 1.3MB). The frequency of sensor cavity and reference cavity was adjusted by three dimples of reference cavity. They are 2mm depth dimple of 2 or 3mm diameter for 10MHz up. Adjustment were within 1MHz, say 0.2~0.5MHz, very good. At first, we do not use heater for temperature stabilization, although heater was tested preliminary. Also two set of additional new electronics were coming today. Next week, we will start tuning of electronics and observing RF signal(714MHz) at Unagi. Cavity mover will also be tested. The optical anchor will be installed late October or November, after precise measurement of temperature. We gave up to measure electrical center offset to mechanical center for these cavities, because of unknown factor of measurement system makes offset very unstable. Actually the recent measurement shows X offset ~50um, Y offset ~150um, although privious measurement was ~150um and ~1500um. We will continue the studies back to with the cold model.

The next meeting will be held in 4 or 25 November, 2004, at 1:30pm-, 3-gokan 325.